OpenStudy (anonymous):

Select the correct answer. (At first I believed it was BCBC but it seems I am wrong)

OpenStudy (anonymous):

(i) Consider the vector A = 3k hat, and any non-zero vector B such that A×B = 0. A)The vector B must be in the x-y plane. B)The vector B must be parallel to k hat. C)The vector B must be parallel to î. D)None of the three statements above is true. (ii) Assume C = 2k hat. Which of the following statements is true? A)If D = (5/2)ĵ, then C×D = 5k hat. B)If D = (−1/2)k hat, then C×D = −1î. C)If D = (−3/2)î, then C×D = 3ĵ. D)None of the three statements above is true. (iii) Consider the vector E = 5î, and any non-zero vector F such that E·F = 0. A)The vector F must be in the x-y plane. B)The vector F must be in the y-z plane. C)The vector F must be in the x-z plane. D)None of the three statements above is true. (iv) Consider the vector G = 5ĵ. Which of the following vectors H satisfies G·H = 1 ? A)H = (1/5)î. B)H = (1/5)î+ĵ+k hat. C)H = î+(1/5)ĵ+k hat. D)H = î+ĵ+(1/5)k hat. E)H = î+ĵ+k hat.